# Dual Prompt Tuning for Domain-Aware Federated Learning

### v1

@@ 5.1 Experimental Setup
Baseline methods. We first consider the baselines of CLIP and its adapted models to federated learning. The Zero-shot CLIP, which infers by aligning images to their class names with a hand-crafted prompt, is a direct baseline to evaluate whether in-domain tuning is necessary for vision-language models in federated learning. We also introduce Single-domain tuning, which applies textual prompt tuning (Zhou et al., 2021) to CLIP only in the local domain, as another baseline to testify whether it is helpful to combine the information across multiple domains. There are also domain-agnostic federated learning approaches based on CLIP such as PromptFL (Guo et al., 2023) and FedCLIP (Lu et al., 2023), which train text prompt and an adapter layer in federated learning fashion, respectively. To further validate the effectiveness of our method, we also compare it with conventional federated learning algorithms FedAvg (McMahan et al., 2017) and FedProx (Li et al., 2020b) that are not based on CLIP. We equip these two baselines by a 50-layer ResNet (He et al., 2016) and a base-scale vision transformer with 16$\times$16 patch size (Dosovitskiy et al., 2021), both being pre-trained on ImageNet-1k (Deng et al., 2009).
Implementation details. For our Fed-DPT, we employ a pre-trained CLIP model with a ViT-Base/16 image encoder, so each textual and visual prompt token has the dimension of 512 and 768, respectively. We set the length of each textual prompt sequence $m=16$ for better robustness, which follows the practice of TPT (Zhou et al., 2021). By default, the number of clients is determined by the number of domains for each dataset, i.e. $n=6$ for DomainNet and $n=4$ for OfficeHome and PACS. We train both our model and the baseline models for 200 epochs and execute the aggregation or broadcast process after every one epoch. We train the ResNet-based models and prompt tokens by a SGD optimizer with 0.01 learning rate, 0.9 momentum, and 0.005 weight decay. Fed-DPT instead uses AdamW (Loshchilov & Hutter, 2019) optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.999$, 5e-4 learning rate, and 0.01 weight decay for transformer-based models. We set the temperature coefficient $\tau_{d}=0.1$ in Equation 5, and set the momentum update ratio $\alpha=0.99$ in Equation 8. If not specified, all reported results are average numbers over three trials.
5.2 Main Results
Table 1 shows the evaluation results of our method and baselines on DomainNet. We observe that our Fed-DPT outperforms the baseline methods by a large margin in terms of average accuracy over six domains. Notably, while the zero-shot CLIP and single-domain tuning protocols fail to obtain reasonable accuracy in the “quickdraw” domain, our Fed-DPT improves this number to 41.6%, which empirically validates the effectiveness of our approach. Benefiting from the technique of prompt learning that introduces very small number of trainable parameters, we find our Fed-DPT to perform very robust to big models. In contrast, conventional methods such as FedAvg (McMahan et al., 2017) and FedProx (Li et al., 2020b) only yield very marginal improvements, or even incur performance degradation when changing the backbone from ResNet-50 (26M parameters) to ViT-Base (86M parameters). Compared to the domain-agnostic prompt learning methods, our Fed-DPT attains higher average accuracy and lower standard deviation (13.8% vs. 16.5% for FedCLIP and 16.4% for PromptFL). This is possibly because our method considers each image’s feature representation in all domains, which makes our predictions more robust to domain shift.
@@ 5.2 Main Results
Z. S. CLIP (Radford et al., 2021) 66.1 40.6 62.3 13.5 80.4 58.5 53.6
@@
FedAvg (ResNet-50 backbone) 40.2 61.1 57.6 33.5 75.6 60.3 54.7
FedAvg (ViT-B/16 backbone) 42.4 60.7 57.0 30.4 79.8 61.1 55.2
FedProx (R-50(Li et al., 2020b) 41.5 62.0 56.8 34.9 79.2 62.6 56.2
FedProx (ViT-B(Li et al., 2020b) 40.5 63.1 57.4 29.7 81.2 59.8 55.3
@@
PromptFL (Guo et al., 2023) 76.0 50.2 70.4 33.5 81.2 67.8 63.2
FedCLIP (Lu et al., 2023) 74.1 48.3 68.5 31.8 80.5 58.6 60.3
Fed-DPT (ours) 77.5 63.1 70.5 41.6 85.7 72.1 68.4
We further evaluate the models on the other two benchmarks and summarize the results in Table 2. The experiments on OfficeHome and PACS also support our conclusion of Fed-DPT’s effectiveness by demonstrating higher average accuracy and lower deviation across domains. Specifically, we improve the zero-shot CLIP by 4.3% average accuracy and 0.3% standard deviation over four domains in OfficeHome. We also observe that overall, the prompt-based methods consistently outperform the conventional federated learning algorithms that require to train the entire model. This confirms the benefits of employing parameter-efficient approaches in federated learning, and explains why we choose to use prompt tuning to address the domain shift issues.
Table 2: Test accuracy (%) on OfficeHome and PACS. Domains include art, clipart, product, and real-world for OfficeHome, and photo, art painting, cartoon, and sketch for PACS. Our results are marked in blue. The best results in each domain are bolded.
@@
5.3 Ablation Studies
Model components. We first dissect our Fed-DPT model to ablate its performance gains. Overall, Fed-DPT comprises two primary components: visual prompts and domain-specific text prompts. By dissecting these components, we get three more variants of our method: 1) Visual Only, it leverages learnable prompt tokens for only image input and uses CLIP’s hand-crafted prompt for texts. 2) Textual Only, it discards the visual prompt tokens of Fed-DPT and uses learnable text prompts only. Note that in the absence of visual prompts, we cannot get the weight $w_{i}$ (see Equation 5 and  6) for each domain, so the text prompts from external clients should also be discarded. We instead aggregate the textual prompts by federated averaging (McMahan et al., 2017). 3) Domain-Agnostic DPT, it retains both Fed-DPT’s visual and textual prompts but decouples them, i.e., we do not perform the weighted sum process in Equation 6, which can be considered as a simple combination of the modes Textual Only and Visual Only.
We summarize the results in Table 3. Since we introduce visual prompt tuning for combining domain information rather than enhancing the visual feature extraction abilities, we do not attach an additional head for the image encoder as in (Jia et al., 2022). Therefore, the Visual Only mode cannot yield significant performance improvements. We also observe that tuning textual prompts results in a 5.5% increase in accuracy, and when tuning them in a federated learning fashion, we achieve an additional 4.1% improvement (Textual Only). Notably, compared to the simple visual-and-textual prompt tuning with 63.5% accuracy, our Fed-DPT achieves a much higher result of 68.4%, which demonstrates the crucial significance of our domain-aware mechanism.
Table 3: Ablation study to model components. We report the average accuracy (%) over six domains in DomainNet. The v. prompt and t. prompt denote whether using visual or textual prompts. Our default setup is marked in blue.
Method federated v. prompt t. prompt domain-specific acc.
@@ 5.3 Ablation Studies
Single-Domain Tuning 59.1 (+5.5)
Visual Only 54.2 (+0.6)
Textual Only 63.2 (+9.6)
Domain-Agnostic DPT 63.5 (+9.9)
Fed-DPT 68.4 (+14.8)
Momentum update, prompt length, and communication frequency. In addition, we consider three more factors that may affect our results. As mentioned in Section 4, we update the external text prompts by exponential moving average to prevent parameters’ sudden change. Here we present comparisons regarding the update mechanism for text prompts in Table 4a, where the accuracy drops by 2.2% in the absence of momentum update. If we train all text prompt tokens in every client, i.e., we disregard the relationship between text prompts and domains, the accuracy drops by 4.4% as it makes Fed-DPT a domain-agnostic approach.
By default, we aggregate the visual prompt tokens by federated averaging, as separately training each token in a specific domain does not yield better performance (see Table 4b). As shown in Table 4c, we set the length of each textual prompt sequence to $m=16$, as it works more robust than a shorter prompt ($m=4$), and when we further increase the length, the model tends to overfit and accuracy drops. In Table 4d we also assess the impact of communication frequency by varying it to 0.5, 1, and 2 training epochs per communication round. It shows that compared to our default setup of one epoch per communication round, more frequent aggregation (0.5 epoch/round) does not lead to improved performance, while conversely, infrequent communication (2 epochs/round) results in a 0.5% accuracy degradation.
@@
Decentralization. By default, we consider each domain in the dataset as a single client, leading to non-identical feature distributions yet the same class distribution across clients. To further testify our method’s effectiveness and flexibility, we conduct a more challenging scenario on DomainNet where each domain is further divided into five clients by Dirichlet sampling, leading to 30 sub-datasets with either non-i.i.d. features or non-i.i.d. categories. Under this setup, we average the text prompt tokens for clients in the same domain at the aggregation step. The results are summarized in Table 5. Compared to our default setting which each domain is considered as one client, our Fed-DPT only has 1.5% accuracy decrease when the dataset is further divided. In contrast, the conventional methods FedAvg and FedProx perform more sensitive to the non-i.i.d categories, with 3.6% and 2.9% accuracy decrease, respectively.
@@
Fed-DPT (ours) 75.8 62.3 69.0 39.5 83.9 70.6 66.9
6 Conclusion
This work introduces Fed-DPT, a novel federated learning approach explicitly designed to address the key challenges of domain shift and communication efficiency. Our method strategically combines CLIP and prompt learning techniques for both visual and textual inputs, thereby enhancing parameter-efficiency and minimizing communication costs, while maintaining robustness in federated optimization involving heterogeneous data. Furthermore, we confront the pervasive issue of domain shift across clients by introducing domain-specific prompts and facilitating correlations between visual and textual representations through self-attention mechanisms. These innovations result in a domain-aware federated learning methodology that consistently demonstrates outstanding effectiveness. Notably, our experiments reveal a remarkable achievement—an average accuracy of 68.4% across six domains in the DomainNet dataset, marking an impressive 14.8% improvement over the original CLIP model. In comparisons with traditional federated learning methods like FedAvg and FedProx, as well as existing domain-agnostic CLIP-based approaches such as PromptFL and FedCLIP, our Fed-DPT consistently outperforms them across three benchmark scenarios.
Appendix
Robustness to few-shot learning. One of the primary advantages of prompt learning is the robustness to few-shot scenarios. We investigate if our dual prompt tuning method retains this merit in the context of federated learning. Therefore, we conduct few-shot learning experiments on DomainNet, employing 1, 2, 4, 8, and 16 training samples per category and per domain. We evaluate the other CLIP-based methods with the same setting, yet only test 16-shot performance for FedAvg as it fails to yield reasonable results with fewer training samples. The corresponding results are summarized in Table 6. As is shown, CLIP-based methods exhibit superior robustness against few-shot learning than FedAvg, which again demonstrates the significant benefits of using parameter-efficient approaches. Also, our Fed-DPT consistently outperforms the baselines in few-shot learning.
Method CLIP-based full 1-shot 2-shot 4-shot 8-shot 16-shot
Single Domain Tuning 59.1 51.1 51.8 53.2 54.7 56.2
FedAvg (ResNet-50) 54.7 - - - - 15.1
FedAvg (ViT-Base/16) 55.2 - - - - 19.7
PromptFL 63.2 51.4 51.8 55.2 57.6 61.2
FedCLIP 60.3 50.8 51.2 52.1 53.4 54.6
Fed-DPT (ours) 68.4 55.4 57.2 60.3 62.7 64.5

### v5

@@ 2 N/A # dng , with
Datasets and Baselines. We evaluate our ADAPT and baseline methods on three domain adaptation image classification benchmarks: the DomainNet (Peng et al. 2019), OfficeHome (Venkateswara et al. 2017), and PACS (Li et al. 2017) datasets, with details can be found in Appendix. We first consider the baselines of CLIP and its adapted models to federated learning. The Zero-shot CLIP, which infers by aligning images to class names with a hand-crafted prompt, is a direct baseline to evaluate whether in-domain tuning is necessary for vision-language models in federated learning. We also introduce Single-domain tuning, which applies textual prompt tuning (Zhou et al. 2021) to CLIP only in the local domain, as another baseline to testify whether it is helpful to combine the information across multiple domains. There are also domain-agnostic federated learning approaches based on CLIP such as PromptFL (Guo et al. 2023), pFedPG (Yang, Wang, and Wang 2023), FedAPT (Su et al. 2024) and FedCLIP (Lu et al. 2023), which train text prompt and an adapter layer in federated learning fashion, respectively. To further validate the effectiveness of our method, we also compare it with conventional federated learning algorithms FedAvg (McMahan et al. 2017) and FedProx (Li et al. 2020) that are not based on CLIP. We equip these two baselines by a 50-layer ResNet (He et al. 2016) and a base-scale vision transformer with 16$\times$16 patch size (Dosovitskiy et al. 2021), both being pre-trained on ImageNet-1k (Deng et al. 2009).
@@
Zero-Shot CLIP (Radford et al. 2021) 66.1 40.6 62.3 13.5 80.4 58.5 53.6
@@
FedAvg (ResNet-50 ) 40.2 61.1 57.6 33.5 75.6 60.3 54.7
FedAvg (ViT-B/16 ) 42.4 60.7 57.0 30.4 79.8 61.1 55.2
FedProx (ResNet-50 (Li et al. 2020) 41.5 62.0 56.8 34.9 79.2 62.6 56.2
FedProx (ViT-B/16(Li et al. 2020) 40.5 63.1 57.4 29.7 81.2 59.8 55.3
@@
PromptFL (Guo et al. 2023) 76.0 50.2 70.4 33.5 81.2 67.8 63.2
FedCLIP (Lu et al. 2023) 74.1 48.3 68.5 31.8 80.5 58.6 60.3
pFedPG (Yang, Wang, and Wang 2023) 73.9 49.2 69.8 32.2 81.4 62.6 61.5
FedAPT (Su et al. 2024) 76.3 49.8 69.2 35.7 81.5 68.2 63.5
ADAPT (ours) 77.5 63.1 70.5 41.6 85.7 72.1 68.4
Implementation details. For our ADAPT, we employ a pre-trained CLIP model with a ViT-Base/16 image encoder, so each textual and visual prompt token has the dimension of 512 and 768, respectively. We set the length of each textual prompt sequence $m=16$ for better robustness, which follows the practice of TPT (Zhou et al. 2021). By default, the number of clients is determined by the number of domains for each dataset, i.e. $n=6$ for DomainNet and $n=4$ for OfficeHome and PACS. We train both our model and the baseline models for 200 epochs and execute the aggregation or broadcast process after every one epoch. We train the ResNet-based models and prompt tokens by a SGD optimizer with 0.01 learning rate, 0.9 momentum, and 0.005 weight decay. ADAPT instead uses AdamW (Loshchilov and Hutter 2019) optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.999$, 5e-4 learning rate, and 0.01 weight decay for transformer-based models. We set the temperature coefficient $\tau_{d}=0.1$ in Equation 4, and set the momentum update ratio $\alpha=0.99$ in Equation 6. If not specified, all reported results are average numbers over three trials.
Main Results
Table 2 shows that our ADAPT significantly outperforms baseline methods on DomainNet, with notably high improvement in the “quickdraw” domain at 41.6% accuracy. This underscores the effectiveness of our prompt learning approach, which requires fewer trainable parameters, enhancing robustness even with larger models. In contrast, traditional methods like FedAvg and FedProx show minimal or negative gains, especially when upgrading from ResNet-50 to ViT-Base. Our ADAPT also achieves higher average accuracy and lower standard deviation compared to domain-agnostic methods, demonstrating better resilience against domain shifts. We further evaluate the models on OfficeHome and PACS, with the results are summarized in Table 3. The experiments on these benchmarks also support our conclusion of ADAPT’s effectiveness by demonstrating higher average accuracy and lower deviation across domains. Specifically, we improve the zero-shot CLIP by 4.3% average accuracy and 0.3% standard deviation over four domains in OfficeHome. We also observe that overall, the prompt-based methods consistently outperform the conventional federated learning algorithms that require to train the entire model. This confirms the benefits of employing parameter-efficient approaches in federated learning, and explains why we choose to use prompt tuning to address the domain shift issues.
Table 3: Test accuracy (%) on OfficeHome and PACS. Domains include art, clipart, product, and real-world for OfficeHome, and photo, art painting, cartoon, and sketch for PACS. Our results are marked in blue. The best results are bolded.
@@
Ablation Studies
We first dissect ADAPT model to ablate its performance gains. ADAPT comprises two primary components: visual prompts and domain-specific text prompts. By dissecting these components, we get three more variants of our method: 1) Visual Only, it leverages learnable prompt tokens for only image input and uses CLIP’s hand-crafted prompt for texts. 2) Textual Only, it discards the visual prompt tokens of ADAPT and uses learnable text prompts only. Note that in the absence of visual prompts, we cannot get the weight $w_{i}$ (see Equation 4 and  5) for each domain, so the text prompts from external clients should also be discarded. We instead aggregate the textual prompts by federated averaging (McMahan et al. 2017). 3) Domain-Agnostic, it retains both ADAPT’s visual and textual prompts but decouples them, i.e., we do not perform the weighted sum process in Equation 5, which can be considered as a simple combination of the modes Textual Only and Visual Only.
We summarize the results in Table 4. Since we introduce visual prompt tuning for combining domain information rather than enhancing the visual feature extraction abilities, we do not attach an additional head for the image encoder as in (Jia et al. 2022). Therefore, the Visual Only mode cannot yield significant performance improvements. We also observe that tuning textual prompts results in a 5.5% increase in accuracy, and when tuning them in a federated learning fashion, we achieve an additional 4.1% improvement (Textual Only). Notably, compared to the simple visual-and-textual prompt tuning with 63.5% accuracy, ADAPT achieves a much higher result of 68.4%, demonstrating the crucial significance of our domain-aware design.
Table 4: Ablation study to model components. We report the average accuracy (%) over six domains in DomainNet. VPT and TPT denote whether using visual or textual prompts.
Method Fed. VPT TPT domain acc.
@@
Single-Domain 59.1
Visual Only 54.2
Textual Only 63.2
Domain-Agnostic 63.5
ADAPT 68.4
Momentum update, prompt length, and communication frequency. We consider three more factors that may affect results. As mentioned in Section Methodology, we update the external text prompts by exponential moving average to prevent parameters’ sudden change. Table 5a presents comparisons regarding the update mechanism for text prompts, where the accuracy drops by 2.2% in the absence of momentum update. If we train all text prompt tokens in every client, i.e., we disregard the relationship between text prompts and domains, the accuracy drops by 4.4% as it makes ADAPT a domain-agnostic approach.
By default, we aggregate the visual prompt tokens by federated averaging, as separately training each token in a specific domain does not yield better performance (see Table 5b). As shown in Table 5c, we set each textual prompt length to $m=16$, as it works more robust than a shorter prompt ($m=4$), and when we further increase the length, the model tends to overfit and accuracy drops. In Table 5d we also assess the impact of communication frequency by varying it to 0.5, 1, and 2 training epochs per communication round. It shows that compared to our default setup of one epoch per communication round, more frequent aggregation (0.5 epoch/round) does not lead to improved performance, while conversely, infrequent communication (2 epochs/round) results in a 0.5% accuracy degradation.
@@
Conclusion
This work introduces ADAPT, a novel federated learning approach explicitly designed to address the key challenges of domain shift and communication efficiency. Our method strategically combines CLIP and prompt learning techniques for both visual and textual inputs, thereby enhancing parameter-efficiency and minimizing communication costs, while maintaining robustness in federated optimization involving heterogeneous data. Furthermore, we confront the pervasive issue of domain shift across clients by introducing domain-specific prompts and facilitating correlations between visual and textual representations through self-attention mechanisms. These innovations result in a domain-aware federated learning methodology that consistently demonstrates outstanding effectiveness. Notably, our experiments reveal a remarkable achievement—an average accuracy of 68.4% across six domains in the DomainNet dataset, marking an impressive 14.8% improvement over the original CLIP model. In comparisons with traditional federated learning methods like FedAvg and FedProx, as well as existing domain-agnostic CLIP-based approaches such as PromptFL and FedCLIP, our ADAPT consistently outperforms them across three benchmark scenarios.
Appendix
Algorithm 1 Training Process of ADAPT
1:Input:
2:      CLIP vision encoder ${\bm{F}}_{V}$, text encoder ${\bm{F}}_{T}$
3:      $n$ local datasets, each ${\bm{D}}_{i}=\left\{(\text{[image]},\text{[class name]})_{j}\right\}_{j=1}^{J}$
4:      Total communication rounds $T$, momentum coefficient $\alpha$
5:Initialization:
6:      Randomly initialize text prompts $[{\bm{P}}_{T}^{1}]^{0},\ldots,[{\bm{P}}_{T}^{n}]^{0}$
7:      Randomly initialize visual prompts $[{\bm{V}}]=\{[{\bm{v}}]_{1},\ldots,[{\bm{v}}]_{n}\}$
8:      Broadcast the pretrained model and prompts to $n$ clients
9:for $t=1$ to $T$ do
10:     # Local training in parallel
11:     for $i=1$ to $n$ do
12:         Keep ${\bm{F}}_{V}$ and ${\bm{F}}_{T}$ frozen
13:         for $j=1$ to $J$ do
14:              Compute ${\bm{f}}_{T}^{k}={\bm{F}}_{T}({\bm{P}}_{T}^{k},\text{[class name]}_{j})$ for $k\in\{1,\ldots,n\}$
15:              Compute ${\bm{f}}_{V}={\bm{F}}_{V}(\text{[cls]},[{\bm{v}}]_{1},\ldots,[{\bm{v}}]_{n},\text{[image]}_{j})$
16:              Extract attention scores ${\bm{w}}=[w_{1},\ldots,w_{n}]$ from ${\bm{F}}_{V}$ using Eq.5
17:              Weighted sum: ${\bm{f}}_{T}=\sum_{k=1}^{n}w_{k}{\bm{f}}_{T}^{k}$
18:              Compute L2 loss: ${\mathcal{L}}=<{\bm{f}}_{V},{\bm{f}}_{T}>/||{\bm{f}}_{V}||\cdot||{\bm{f}}_{T}||$
19:              Update $[{\bm{v}}]_{1},\ldots,[{\bm{v}}]_{n}$ and ${\bm{P}}_{T}^{i}$ by ${\mathcal{L}}$
20:              Update ${\bm{P}}_{T}^{k},k\in\{1,\ldots,n\},k\neq i$ by momentum: ${\bm{P}}_{T}^{k}=\alpha{\bm{P}}_{T}^{k}+(1-\alpha)[{\bm{P}}_{T}^{k}]^{t-1}$
21:         end for
22:     end for
23:     # Global aggregation in the server
24:     Average $[{\bm{V}}]=\frac{1}{n}\sum_{k=1}^{n}[{\bm{V}}]^{k}$, where $[{\bm{V}}]^{k}=\{[{\bm{v}}]_{1},\ldots,[{\bm{v}}]_{n}\}$ obtained from #$k$ client
25:     Assign $[{\bm{P}}_{T}^{k}]^{t}={\bm{P}}_{T}^{k}$, where ${\bm{P}}_{T}^{k}$ obtained from #$k$ client
26:     Broadcast $[{\bm{V}}],[{\bm{P}}_{T}^{k}]^{t}(k\in\{1,\ldots,n\})$ to all clients
27:end for
Datasets. We evaluate our ADAPT and baseline methods on the following three domain adaptation image classification benchmarks:
  • DomainNet (Peng et al. 2019). The DomainNet dataset has around 600,000 images spanning 345 categories from six domains, which covers diverse image styles including clipart, infograph, painting, quickdraw, real, and sketch.
  • OfficeHome (Venkateswara et al. 2017). The OfficeHome dataset consists of approximately 15,500 images depicting everyday objects in 65 classes. It further categorizes the images into four domains: art, clipart, product, and real-world.
  • PACS (Li et al. 2017). The PACS dataset contains around 10,000 images drawn from seven categories and four domains, including photo, sketch, cartoon, and painting styles.
Communication Costs. ADAPT markedly reduces communication overhead in federated learning by only transferring domain prompts, contrary to standard methods that share all trainable parameters. To provide a clear comparison, we have included the following results in Table 6 and Figure 4. An additional benefit of this approach is its ability to produce favorable results without requiring a substantial volume of training data. As shown in Table 8(in the supplementary material), we obtain very competitive few-shot results by our prompt tuning technique. In practice, We avoid fine-tuning the CLIP model to maintain its visual-language alignment. Fine-tuning large models such as CLIP escalates communication expenses and impedes the rate of convergence. With an equivalent number of training iterations, the fine-tuning protocol often falls short to prompt learning.
Table 6: Comparison of Parameters and accuracy (%) to fine-tuning protocols on DomainNet dataset. Our results are marked in blue. The best results in each domain are bolded
Method Learnable params acc.
FedAvg (McMahan et al. 2017) 86M 57.6
FedProx (Li et al. 2020) 86M 58.1
ADAPT (ours) 16.9k 68.4
Figure 4: Comparison of performance to fine-tuning protocols on DomainNet dataset
Method CLIP-based full 1-shot 2-shot 4-shot 8-shot 16-shot
Single Domain Tuning 59.1 51.1 51.8 53.2 54.7 56.2
FedAvg (ResNet-50) 54.7 - - - - 15.1
FedAvg (ViT-Base/16) 55.2 - - - - 19.7
PromptFL 63.2 51.4 51.8 55.2 57.6 61.2
FedCLIP 60.3 50.8 51.2 52.1 53.4 54.6
ADAPT (ours) 68.4 55.4 57.2 60.3 62.7 64.5
@@ Conclusion
ADAPT (ours) 75.8 62.3 69.0 39.5 83.9 70.6 66.9
Robustness to few-shot learning. One of the primary advantages of prompt learning is the robustness to few-shot scenarios. We investigate if our dual prompt tuning method retains this merit in the context of federated learning. Therefore, we conduct few-shot learning experiments on DomainNet, employing 1, 2, 4, 8, and 16 training samples per category and per domain. We evaluate the other CLIP-based methods with the same setting, yet only test 16-shot performance for FedAvg as it fails to yield reasonable results with fewer training samples. The corresponding results are summarized in Table 7. As is shown, CLIP-based methods exhibit superior robustness against few-shot learning than FedAvg, which again demonstrates the significant benefits of using parameter-efficient approaches. Also, our ADAPT consistently outperforms the baselines in few-shot learning.
Decentralization. By default, we consider each domain in the dataset as a single client, leading to non-identical feature distributions yet the same class distribution across clients. To further testify our method’s effectiveness and flexibility, we conduct a more challenging scenario on DomainNet where each domain is further divided into five clients by Dirichlet sampling, leading to 30 sub-datasets with either non-i.i.d. features or non-i.i.d. categories. Under this setup, we average the text prompt tokens for clients in the same domain at the aggregation step. The results are summarized in Table 8. Compared to our default setting which each domain is considered as one client, our ADAPT only has 1.5% accuracy decrease when the dataset is further divided. In contrast, the conventional methods FedAvg and FedProx perform more sensitive to the non-i.i.d categories, with 3.6% and 2.9% accuracy decrease, respectively.